# One Stone with Two Birds: A Null-Text-Null Frequency-Aware Diffusion Models for Text-Guided Image Inpainting

### v2

@@ 2.3.1 Null-Text Low-Frequency Aware Denoising Process
Formally, we initialize the denoising process with random Gaussian noise $z^{un}_{T}\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{8}}$ and null-text $c_{\emptyset}$ as the input to the denoising network $\epsilon_{\theta}(z_{T}^{un},T,\tau_{\theta}(c_{\emptyset}))$ for null-text denoising process. To preserve the unmasked regions, at each step $t$, we replace the unmasked regions of denoising result $z^{un}_{t}$ with those from the diffusion result $z^{gt}_{T-t}$ at the (TtT-t)-th step, as per the following:
@@ 2.3.2 Text-Guided Denoising Process
To inpaint the masked regions as per text prompts, we resort to the text-guided denoising process, together with low-frequency band substitution from null-text denoising process (Sec.2.3.1). Specifically, we initialize the denoising process with random Gaussian noise $z^{text}_{T}\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{8}}$ and text prompts $c$ as the input to the denoising network $\epsilon_{\theta}(z_{T}^{text},T,\tau_{\theta}(c))$ for text-guided denoising process for masked regions. As inspired by Fig.2, we preserve the unmasked regions as per their low-frequency band, since, instead of mid-frequency or even high-frequency band, such information can be easily modulated by text prompt. To this end, we propose a plug-and-play Denoised Low-Frequency Band layer (Fig.4(a)) to substitute the latent representation $z_{t}^{text}$ within the text-guided denoising process, by ${z}^{un}_{t}$ from the null-text denoising process (see Sec.2.3.1) for the $t$-th step, which is formulated as:
@@ 2.3.3 Null-Text Mid-Frequency Aware Denoising Process
We resort to another null-text denoising process guided by the denoised mid-frequency band from the second text-guided denoising process (Sec.2.3.2) for the $t$-th step of the early stage. Particularly, we initialize the denoising process with random Gaussian noise $z^{mid}_{T}\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{8}}$ and null-text $c_{\emptyset}$ as input to the denoising network $\epsilon_{\theta}(z_{T}^{mid},T,\tau_{\theta}(c_{\emptyset}))$. We further exploit the denoised mid-frequency result from $\tilde{z}_{t}^{text}$ , to substitute the denoised mid-frequency band ${z}^{mid}_{t}$ from the last null-text denoising process, throughout a plug-and-play Denoised Mid-Frequency Band layer (as illustrated in Fig.4(b)), which is formulated below:
$\begin{split}\tilde{z}^{mid}_{t}=&\,\text{IDCT}\left(\text{DCT}(\tilde{z}_{t}^{text})\odot m_{\text{mid}}+\text{DCT}({z}^{mid}_{t})\odot(1-m_{\text{mid}})\right),\end{split}$ (8)
@@ 2.4 Late Stage of Text-Guided Denoising Process
Following the discussion above, for final late stage of denoising process, we perform the text-guided denoising process based on the early null-text mid-frequency guided denoising process (Sec.2.3.3), the low-frequency band within the masked regions is denoised throughout the path to text prompts alignment, together with the stable mid-frequency band (see Fig.2(b)) from the early text-guided denoising process (Sec.2.3.2), the masked regions are well inpainted. However, the low-frequency band for unmasked regions biases a lot from the ground truth from the early null-text denoising process (Sec.2.3.3), and augmented by the influence from the text prompts for late stage. To this end, we substitute the unmasked regions from early stage of the diffusion process into the denoising results $z^{mid}_{t}$ for the $t$-th step during the late stage, which is formulated as:
@@ 3.2 Comparison with State-of-the-arts
Quantitative and Qualitative analysis. The quantitative results in Table.1 summarizes our findings below: NTN-Diff enjoys larger IR, HPS v2, PSNR and CLIP Score, together with smaller MSE and LPIPS than the competitors. Notably, BrushNet remains the large performance margins (at most 0.73% for IR, 0.62% for HPS, 8.59% for PSNR, 30.18% for MSE, 15.51% for LPIPS and 0.23% for CLIP Score) compared to NTN-Diff in Table.1. We also present the quantitative results of NTN-Diff with the pixel-level blending operation of ju2024brushnet , named NTN-Diff*, to preserve the unmasked regions, which demonstrates the ability to tame the hybrid frequency issue, the results further verifies the intuition in Sec.1 – NTN-Diff achieves the semantics consistency between mid-and-low frequency bands across masked and unmasked regions, while preserving unmasked regions. We illustrate the above intuitions in Fig.5. (See more analysis of computational efficiency in Sec.D of the Appendix, along with more compared results for Editbench wang2023imagen can be seen in Sec.E of the Appendix)
@@ 3.3 Ablation Studies
3.3.1 Discussion on the Different Denoising Process of the Early Stage
@@ 4 Conclusion
Acknowledgments This research is supported by National Natural Science Foundation of China (U21A20470, 62172136, 72188101); Institute of Advanced Medicine and Frontier Technology (2023IHM01080); CCF-NetEase ThunderFire Innovation Research Funding (CCF-Netease 202513).

### v4

@@ 2.3.1 Null-Text Low-Frequency Aware Denoising Process
Formally, we initialize the denoising process with random Gaussian noise $z^{un}_{T}\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{8}}$ and null-text $c_{\emptyset}$ as the input to the denoising network $\epsilon_{\theta}(z_{T}^{un},T,\tau_{\theta}(c_{\emptyset}))$ for null-text denoising process. To preserve the unmasked regions, at each step $t$, we replace the unmasked regions of denoising result $z^{un}_{t}$ with those from the diffusion result $z^{gt}_{T-t}$ at the (TtT-t)-th step to yield $\hat{z}^{un}_{t}$, as per the following:
@@ 2.3.2 Text-Guided Denoising Process
To inpaint the masked regions as per text prompts, we resort to the text-guided denoising process, together with low-frequency band substitution from null-text denoising process (Sec.2.3.1). Specifically, we initialize the denoising process with random Gaussian noise $z^{text}_{T}\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{8}}$ and text prompts $c$ as the input to the denoising network $\epsilon_{\theta}(z_{T}^{text},T,\tau_{\theta}(c))$ for text-guided denoising process for masked regions. As inspired by Fig.2, we preserve the unmasked regions as per their low-frequency band, since, instead of mid-frequency or even high-frequency band, such information can be easily modulated by text prompt. To this end, we propose a plug-and-play Denoised Low-Frequency Band layer (Fig.4(a)) to substitute the latent representation $z_{t}^{text}$ within the text-guided denoising process, by ${z}^{un}_{t}$ from the null-text denoising process (see Sec.2.3.1) for the $t$-th step to yield $\tilde{z}_{t}^{text}$, which is formulated as:
@@ 2.3.3 Null-Text Mid-Frequency Aware Denoising Process
We resort to another null-text denoising process guided by the denoised mid-frequency band from the second text-guided denoising process (Sec.2.3.2) for the $t$-th step of the early stage. Particularly, we initialize the denoising process with random Gaussian noise $z^{mid}_{T}\in\mathbb{R}^{4\times\frac{H}{8}\times\frac{W}{8}}$ and null-text $c_{\emptyset}$ as input to the denoising network $\epsilon_{\theta}(z_{T}^{mid},T,\tau_{\theta}(c_{\emptyset}))$. We further exploit the denoised mid-frequency result from ${z}_{t}^{text}$ , to substitute the denoised mid-frequency band ${z}^{mid}_{t}$ from the last null-text denoising process, throughout a plug-and-play Denoised Mid-Frequency Band layer (as illustrated in Fig.4(b)) to yield $\tilde{z}^{mid}_{t}$, which is formulated below:
$\begin{split}\tilde{z}^{mid}_{t}=&\,\text{IDCT}\left(\text{DCT}({z}_{t}^{text})\odot m_{\text{mid}}+\text{DCT}({z}^{mid}_{t})\odot(1-m_{\text{mid}})\right),\end{split}$ (8)
@@ 2.4 Late Stage of Text-Guided Denoising Process
Following the discussion above, for final late stage of denoising process, we perform the text-guided denoising process based on the early null-text mid-frequency guided denoising process (Sec.2.3.3), the low-frequency band within the masked regions is denoised throughout the path to text prompts alignment, together with the stable mid-frequency band (see Fig.2(b)) from the early text-guided denoising process (Sec.2.3.2), the masked regions are well inpainted. However, the low-frequency band for unmasked regions biases a lot from the ground truth from the early null-text denoising process (Sec.2.3.3), and augmented by the influence from the text prompts for late stage. To this end, we substitute the unmasked regions from early stage of the diffusion process into the denoising results $z^{mid}_{t}$ for the $t$-th step during the late stage to yield $\hat{z}^{mid}_{t}$, which is formulated as:
@@ 3.2 Comparison with State-of-the-arts
Quantitative and Qualitative analysis. The quantitative results in Table.1 summarize our findings below: NTN-Diff enjoys larger IR, HPS v2, PSNR and CLIP Score, together with smaller MSE and LPIPS than the competitors. Notably, BrushNet remains the large performance margins (at most 0.73% for IR, 0.62% for HPS, 8.59% for PSNR, 30.18% for MSE, 15.51% for LPIPS and 0.23% for CLIP Score) compared to NTN-Diff in Table.1. We also present the quantitative results of NTN-Diff with the pixel-level blending operation of ju2024brushnet , named NTN-Diff*, to preserve the unmasked regions, which demonstrates the ability to tame the hybrid frequency issue, the results further verifies the intuition in Sec.1 – NTN-Diff achieves the semantics consistency between mid-and-low frequency bands across masked and unmasked regions, while preserving unmasked regions. We illustrate the above intuitions in Fig.5. (See more analysis of computational efficiency in Sec.D of the Appendix, along with more compared results for Editbench wang2023imagen can be seen in Sec.E of the Appendix)
@@ 3.3 Ablation Studies
3.3.1 Discussion on Different Denoising Processes of the Early Stage
@@ 4 Conclusion
Acknowledgments This research is supported by National Natural Science Foundation of China (U21A20470, 62172136, 72188101); Institute of Advanced Medicine and Frontier Technology (2023IHM01080); The computation is completed on the HPC Platform of Hefei University of Technology.